OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(2c) = √5^(4c – 10) ? c = 4 c = – 5 / 2 c = 7 c = 7 / 2 Which of the following is the solution to the equation 25^(2c) = √5^(4c – 10) ? c = 4 c = – 5 / 2 c = 7 c = 7 / 2 @Mathematics

OpenStudy (anonymous):

\[y=25^{2c}<=>\ln y=2c \ln 25\] \[y=\sqrt{5}^{(4c-10)} <=> \ln y=(4c-10)\ln (\sqrt{5})\] So \[2c \ln 25=(4c-10)\ln (\sqrt{5})\] \[2c \ln (5^2)=(4c-10)\ln (5^{0.5})\] \[4c\ln5=\frac{1}{2}(4c-10)\ln5\] \[4c=2c-5\] \[c=-\frac{5}{2}\]